OpenStudy (anonymous):

graph slope -1/5 y-intercept (0,2)

OpenStudy (trojanpoem):

The equation of a line is y = mx + b. m = slope = -1/5 , b = 2 y = -1/5 x + 2

OpenStudy (trojanpoem):

We know that the line intercepts y-axis at 2 so we will start by this point.

OpenStudy (anonymous):

so do i start at 2,0 on the y axis?

OpenStudy (trojanpoem):

Yeah, and sorry for slow writing I am lagging.

OpenStudy (anonymous):

how do i get the -1/5?

OpenStudy (trojanpoem):

From this point go down 1 in y axis and 5 in x-axis

OpenStudy (trojanpoem):